OpenStudy (anonymous):

Find the point(s) on the graph of y=(2x^4+1)(x-5) where the slope is one. I think I started on the right path not sure. I took the derivative of y and set it to zero???

OpenStudy (noelgreco):

The value of the first derivative IS the slope of the tangent line so set y' = 1.

OpenStudy (anonymous):

I meant set it equal to 1

OpenStudy (anonymous):

Yeah that's what I did. I just have trouble solving it. I got 10x^4-40x^3+x-5=0. I was like how do I solve that. If you can help I would really appreciate it

OpenStudy (noelgreco):

You've got to take the derivative of that expression and set it = to one.

OpenStudy (noelgreco):

oops. the derivative is actually 10x^4 - 40 x^3 +1

OpenStudy (anonymous):

I did take the derivative of the expression and set it equal to one. This is my work: y=(2x^4+1)(x-5) y'=(2x^4+1)(1)+(x-5)(8x^3+1) simplify further and you get y'= 10x^4-40x^3+x-4 then i set y'=1 and it gave me: 10x^4-40x^3+x-5=0

OpenStudy (anonymous):

Maybe I'm not doing it right

OpenStudy (noelgreco):

You have the derivative of (2x^4 + 1) as 8x^3 + 1. The 1 is a constant. The derivative should simply be 8x^3.

OpenStudy (anonymous):

ahhh. That's why it is always good to check and recheck your work. Thanks. I'm going to rework it and see what happens

OpenStudy (anonymous):

ok so now I got 10x^4-40x^3+1=1

OpenStudy (anonymous):

Bring one to the other side to get 10x^4-40x^3=0

OpenStudy (anonymous):

I know you can take out a 10x^3 to get:10x^3(x-4)=0. So how would I go about solving this please. Is the quadratic formula feasible? I didn't think it was

OpenStudy (noelgreco):

\[10x ^{3} (x-4)=0\] That gives you rwo locations where the slope of the tangent line = 1.

OpenStudy (anonymous):

so do U solve for 10x^2 and x-4?

OpenStudy (anonymous):

How did you get two locations? I got only one. x=4

OpenStudy (anonymous):

My points are (4,-513). Don't know if it is right, but that's what I am sticking with as of right now

OpenStudy (noelgreco):

The other point is x=0 (0,-5)